# PortraitTalk: Towards Customizable One-Shot Audio-to-Talking Face Generation

### v1

@@ Abstract
Audio-driven talking face generation is a challenging task in digital communication. Despite significant progress in the area, most existing methods concentrate on audio-lip synchronization, often overlooking aspects such as visual quality, customization, and generalization that are crucial to producing realistic talking faces. To address these limitations, we introduce a novel, customizable one-shot audio-driven talking face generation framework, named PortraitTalk. Our proposed method utilizes a latent diffusion framework consisting of two main components: IdentityNet and AnimateNet. IdentityNet is designed to preserve identity features consistently across the generated video frames, while AnimateNet aims to enhance temporal coherence and motion consistency. This framework also integrates an audio input with the reference images, thereby reducing the reliance on reference-style videos prevalent in existing approaches. A key innovation of PortraitTalk is the incorporation of text prompts through decoupled cross-attention mechanisms, which significantly expands creative control over the generated videos. Through extensive experiments, including a newly developed evaluation metric, our model demonstrates superior performance over the state-of-the-art methods, setting a new standard for the generation of customizable realistic talking faces suitable for real-world applications.
Introduction
In recent years, audio-to-talking face generation has gained significant attention within the artificial intelligence community, primarily due to its broad spectrum of applications. These applications span across video content creation, animation production, video dubbing, and immersive experiences in the metaverse. The primary objective of audio-to-talking face generation is to create realistic talking face videos that are precisely synchronized with the provided audio input (Nazarieh et al. 2024; Tan et al. 2024a).
Figure 1: Talking faces generated by PortraitTalk.
The studies in this area can be classified into three main categories: enhancing visual quality, improving audio-lip synchronization, and incorporating emotion into generated talking faces. While previous methodologies have shown promising results, they often produce videos that lack control of head poses, emotional expressions, identity consistency, and facial detail customization. An intuitive approach to overcoming these challenges involves processing individual aspects of facial motion (e.g., mouth movements, head orientation, and eye blink) separately, and integrating them during the rendering phase (Liang et al. 2022; Chen et al. 2019a; Zhou et al. 2020). However, this strategy presents significant drawbacks. Primarily, many of these methods depend heavily on predefined pose coefficients derived from 3D face models (Tan, Ji, and Pan 2024a; Peng et al. 2024; Xu et al. 2024; Peng et al. 2023), which increases the requirement of datasets and introduces potential inaccuracies (Tan et al. 2024a). In addition, these techniques require the entire network to be retrained from scratch when introducing a new identity.
Current methods exhibit limited applicability in multimodal contexts due to their reliance solely on audio (audio-driven talking face generation (Zhou et al. 2020; Gan et al. 2023; Prajwal et al. 2020)) or video (video-driven talking face generation (Ji et al. 2022; Pang et al. 2023; Gong et al. 2023)). Recently, text-driven talking face generation has attracted increasing attention (Song et al. 2022; Li et al. 2021; Tan, Ji, and Pan 2024b; Jang et al. 2024). However, it remains significantly under-explored and is frequently overlooked in existing frameworks. Consequently, there is a pressing need for a comprehensive framework capable of generating talking faces that are controllable across multiple modalities. This capability is critically important for enabling more versatile and realistic results in various domains, including virtual assistants and animated characters. This framework is expected to simultaneously ensure precise control over facial details, achieve accurate audio-lip synchronization and manage background and environment. In this paper, we introduce a customizable one-shot audio-to-talking face generation framework (PortraitTalk) designed to address the above limitations with the assistance of decoupled cross-attention mechanism (Figure 1).
PortraitTalk has two pivotal components: IdentityNet and AnimateNet. IdentityNet preserves consistent identity across video frames, while integrating semantics from text descriptions into the generated faces. To enable customization, we integrate text and image embeddings in IdentityNet through decoupled cross-attention (Ye et al. 2023a). Traditional approaches typically concatenate text-image pairs in text-driven talking face generation (Li et al. 2021; Ma et al. 2023) or audio-image pairs in audio-driven talking face generation (Shen et al. 2023; Ji et al. 2022; Shen et al. 2024) within a singular attention block. We argue that this approach restricts generative models from thoroughly learning and integrating identity-specific facial details and textual descriptions. Through extensive experiments, our findings demonstrate that PortraitTalk significantly enhances high-fidelity identity preservation, consistent talking-face generation, and overall controllability. Other critical quality factors in talking face generation are audio-lip synchronization and temporal consistency. Even minor artifacts or head displacements can substantially degrade the visual experience. AnimateNet leverages a pre-trained motion generator (Ye et al. 2023b) to map audio inputs to their corresponding facial landmarks effectively. This mapping is enhanced by head placement guidance, ensuring that the visual output remains coherent and synchronized throughout the generated video. By tackling these challenges, PortraitTalk represents a significant step forward in audio-to-talking face generation, offering a more robust and flexible solution that enhances the realism and practicality of generated talking face videos.
@@ Introduction
Despite the notable progress in audio-to-talking face generation, the existing evaluation metrics often concentrate solely on isolated attributes such as visual quality or lip synchronization. Consequently, these metrics cannot provide a comprehensive assessment that reflects the holistic performance of a method. There exists an imperative need for a new evaluation metric that encompasses both spatial and temporal consistencies to facilitate thorough evaluation. Spatial consistency is crucial as it ensures the coherence and stability of the generated facial frames, thereby preserving the integrity of identity and expression throughout the video. Temporal consistency, on the other hand, assesses the smoothness and natural flow of the transition over time, which are vital for producing lifelike and persuasive talking faces. In this paper, we propose a novel Audio-Driven Facial Dynamics (ADFD) measurement. ADFD addresses the above deficiencies by integrating both spatial and temporal factors, thereby providing a reliable measure of the overall quality of a generated talking face video.
The main contributions of PortraitTalk include:
@@
  • To the best of our knowledge, this is the first customizable one-shot audio-to-talking face generation method. We use three different modalities in the proposed PortraitTalk framework.
@@
  • We propose a new evaluation metric that effectively measures video quality in both spatial and temporal aspects.
@@
  • Qualitative and quantitative experiments demonstrate the effectiveness of PortraitTalk in enhancing synchronization, realism, and customization of generated talking face videos as compared with the state-of-the-art methods.
Related Work
Audio-driven Talking Face Generation
@@
In this field, various methodologies have been proposed, with intermediately guided talking face generation being particularly common. Such models leverage intermediate representations, such as 2D facial landmarks or 3D faces, to effectively bridge the gap between audio input and visual output. These methods (Zhou et al. 2020, 2021; Gan et al. 2023) typically consist of two main components: the first predicts intermediate representations from audio, and the second uses these features to generate talking faces. This two-stage methodology facilitates control over the synchronization between audio and visual data, thereby enhancing the realism and accuracy of the generated talking faces.
Figure 2: An overview of PortraitTalk that has two main components: IdentityNet and AnimateNet. Text and identity embeddings are derived from pre-trained text and face encoders, with a pre-trained projection layer mapping identity features into the text embedding dimension. These features are integrated into IdentityNet using a decoupled cross-attention mechanism to capture subtle facial characteristics better. Simultaneously, facial motions corresponding to the input speech, enhanced by head placement guidance, are processed through AnimateNet to ensure dynamic and temporal coherence. In our framework, a latent diffusion model serves as the foundational rendering mechanism. The structural attention block incorporates head placement guidance and facial landmark mapping. For simplicity, these elements are represented within a single block.
@@
Chen et al. (2019a) proposed a method that converts audio into mouth landmarks to generate talking faces. However, as it only predicts movements for the lower half of the face and uses the upper half from ground truth frames, the resulting videos often lack realism. Zhou et al. (2020) introduced MakeItTalk, a one-shot audio-to-talking-face generation model based on facial landmarks. While it can handle various identities, MakeItTalk struggles with precise audio-lip synchronization and often fails to deliver high-quality talking faces. The Audio2Head (Wang et al. 2021) model generates talking faces by extracting relatively dense motion fields from audio, but the faces often exhibit distortions and lack consistency in identity preservation. SadTalker (Zhang et al. 2023) has significantly enhanced the generalization capability through a learned latent space. Despite this advancement, the head movements, which are generated based on predefined motion coefficients, frequently lack realism. IP-LAP (Zhong et al. 2023), a transformer-based generative model, utilizes audio inputs and sketches to produce talking faces. Although it achieves stable head movements, it falls short in synchronizing lip movements with audio, particularly with identities unseen during training.
Emotional Talking Face Generation
@@
Emotional talking face generation has attracted increasing attention, primarily due to its potential to enhance communication and its utility in the entertainment industry. Traditional methods have predominantly used discrete emotion labels to model expressions. For example, Goyal et al. (2023) developed a model inspired by Wav2Lip (Prajwal et al. 2020), integrating an emotion label encoder, emotion-specific loss function, and an emotion discriminator. This model, however, primarily modifies the mouth region, thereby producing videos that lack realistic appearance variations. Furthermore, it exhibits limited effectiveness in displaying diverse expressions and fails to maintain an identity consistency across the generated frames.
@@
Label-based emotional talking face generation typically struggles to produce controllable and fine expressions. To address these limitations, recent developments advocate the use of emotional videos as references (Sun et al. 2024; Tan et al. 2024b; Ji et al. 2022; Tan et al. 2024a). This approach allows for the injection of expressions directly derived from other videos, enhancing the expressiveness of generated faces. For example, EAMM (Ji et al. 2022) uses an emotional source video to generate emotional talking faces. Nonetheless, this method often struggles with identity consistency, exhibits noticeable irregularities, and encounters audio-lip synchronization issues. These challenges underscore the necessity for the development of a model that can integrate emotion with audio-visual semantics more effectively. Similarly, EDTalk (Tan et al. 2024a) employs emotions from reference-style videos to enhance the expressiveness of talking faces. While this method substantially enhances expressiveness, it sometimes does so at the cost of visual quality.
The Proposed Method
@@
PortraitTalk is a one-shot audio-driven talking face generation framework with aligned audio-lip movements by using only reference identity images and the corresponding speech. Figure 2 illustrates the pipeline of PortraitTalk. We also enhance the design by integrating text prompt embedding, which enables PortraitTalk to edit video attributes including style, expression, and appearance customization. PortraitTalk has two essential components. First, IdentityNet efficiently embeds high-level and low-level semantics from text and image prompts, ensuring the preservation of precise facial details across the generated video frames. Second, AnimateNet employs condition control features along with temporal consistency mechanisms to facilitate smooth transitions within the video.
IdentityNet
The proposed IdentityNet adopts the architecture of Stable Diffusion (Rombach et al. 2022) (SD1.5) and serves as the primary mechanism for preserving the facial characteristics throughout a generated video. The backbone network retains the original weights of the UNet in SD, enabling it to effectively maintain the generation capability of SD. This provides a more robust initialization for PortraitTalk.
@@
To reduce the inconsistency among synthesized frames and enhance the reconstruction capability of IdentityNet, the model undergoes finetuning with masked images. This strategy involves random corruption of images and subsequent reconstruction of the missing segments. By finetuning IdentityNet using masked images, the model’s focus shifts towards the global characteristics of images, such as the overall head shape and style associated with a specific identity (Chen et al. 2024). The fine-tuning process follows the training strategy of stable diffusion. Randomly masked frames are encoded by passing through the encoder to obtain a masked latent representation $\mathbf{z}_{t}^{m}$. Subsequently, a forward and backward diffusion process is applied across $T$ time steps within this masked latent space. The denoising U-Net ($\epsilon_{\theta}$) is trained to reconstruct the original image by estimating the amount of noise ($\epsilon$) introduced at each forward step. This is achieved by the mask reconstruction loss:
@@
$L^{mask}=\mathbb{E}_{t,\mathbf{z}_{t},c,\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t}^{m},t,\mathbf{c})\|^{2}\right]$ (1)
@@
where $t$ is a sampled time step, $z_{t}^{m}$ is the noisy masked latent at time step $t$, and $c$ is the condition set (text and image). This loss function guides the network towards accurate noise estimation and effective image restoration.
@@
Despite the effectiveness of SD in image generation, it struggles to maintain consistent identity characteristics across consecutive frames. Moreover, given our model’s goal of enabling customization during the generation process, it is essential to incorporate user prompts into the denoising blocks effectively. This ensures that the generated talking faces consistently exhibit facial details that align with the input text and reference image prompts. To tackle these challenges, we employ a decoupled cross-attention mechanism that processes text and image features in parallel, allowing for more targeted and coherent feature integration. Specifically, first, text and reference image features are extracted. Then the features are processed by a pre-trained decoupled cross attention (Ye et al. 2023a) module and embedded into the IndentityNet backbone.
@@
In preceding studies (Ye et al. 2023a; Li et al. 2024), CLIP has been widely used as an image encoder. However, it has proven inadequate for extracting strong facial features in audio-to-talking face generation. This deficiency often results in unstable facial representations across frames, while barely reflecting text prompt guidance in generated video. CLIP, a contrastive language–image model, is primarily trained on text-image pairs from the natural image domain. This training predisposes CLIP to prioritize broad and ambiguous features such as color and style, which are less effective for detailed face image understanding (Wang et al. 2024). Such limitations pose significant challenges for our framework. This task necessitates precise identity preservation, wherein robust semantic representation of facial features is crucial. To achieve this, we initially eliminate background elements from reference images and subsequently extract facial embeddings utilizing a pre-trained face recognition model (Wang et al. 2024). This approach has demonstrated enhanced facial fidelity across generated frames. Simultaneously, text prompts are processed using the CLIP text encoder and facial embeddings are mapped to the same dimensional space as text embeddings via a pre-trained projection layer (Ye et al. 2023a). Both text and facial representations are then integrated into the denoising UNet using decoupled cross-attention blocks, enhancing the facial fidelity and editability of the generated talking faces.

### v2

# PortraitTalk: Towards Robust and Customizable Audio-to-Talking Face Generation

@@ Abstract
Recent advancements in audio-driven talking face generation have largely focused on achieving accurate audio-lip synchronization, often at the expense of visual realism, personalization, and generalizability, which are key factors in producing convincing talking face videos. To address these limitations, we present PortraitTalk, a robust and customizable audio-driven talking face generation framework designed for enhanced visual quality and user control. Our method is based on a latent diffusion architecture comprising two key modules: IdentityNet and AnimateNet. Specifically, IdentityNet preserves the identity across the generated video frames, and AnimateNet ensures smooth and coherent facial motion over time. This approach reduces reliance on reference-style videos prevalent in existing approaches. Another notable contribution of our work is the integration of text prompts through decoupled cross-attention mechanisms, enabling greater creative flexibility in video generation. Last, we propose a new evaluation metric, namely Audio-Driven Facial Dynamics (ADFD), that effectively measures the quality of a generated video in terms of both spatial and temporal aspects. Extensive experiments demonstrate that our method outperforms the state-of-the-art approaches, offering a robust and practical solution for realistic, customizable talking face synthesis driven by audio input.
I Introduction
Recently, audio-to-talking face generation has gained significant attention within the generative AI community, primarily due to its broad spectrum of applications, including video content creation, animation production, video dubbing, etc. The aim of audio-to-talking face generation is to create realistic talking face videos precisely synchronized with the provided audio input [1, 2]. The studies in this area can be broadly divided into three categories: enhancing visual quality, improving audio–lip synchronization, and incorporating emotion into generated talking faces. While existing methods have achieved notable progress in each aspect, they often struggle to jointly preserve head pose control, expressive variation, identity consistency, and fine-grained facial detail customization. One common strategy to address these limitations is to decompose facial motion into separate components, such as mouth movements, head orientation, and eye blinks, and then integrate them during the rendering phase [3, 4, 5]. Although this modular approach can improve controllability, it is not without difficulties. Many implementations rely heavily on predefined pose coefficients extracted from 3D face models [6, 7, 8, 9, 10], which increases dataset requirements and can propagate modeling inaccuracies [2]. Moreover, such systems typically require full retraining when adapting to a new identity, which limits their scalability in practical applications.
In addition to these motion modeling constraints, most existing methods also lack flexibility in handling multiple input modalities for richer control. Current talking face generation frameworks are predominantly driven by either audio [5, 11, 12] or video [13, 14, 15] inputs, offering limited adaptability in multimodal contexts. Recently, text-driven talking face generation has attracted growing interest [16, 17, 18, 19, 20]. However, its potential for fine-grained customization remains largely underexplored. Existing approaches often employ textual descriptions in a narrow way, primarily for reconstructing or setting the scene, rather than enabling rich control over visual attributes. In practice, a more capable framework should harness textual input to manipulate a wide range of elements, from facial attributes and expressions to background composition, object presence, and overall visual style. Such flexibility is essential for producing versatile and realistic results in applications spanning virtual assistants, animated characters, and personalized digital content. Beyond customization, practical deployment also demands the ability to generalize to identities unseen during training and to produce realistic results from only a few reference images. This capability removes the need for costly identity-specific retraining and extends applicability to real-world scenarios where visual data is scarce. Achieving this while supporting multimodal control remains challenging, requiring precise identity preservation, temporal coherence, and accurate audio–lip synchronization.
To mitigate the above issues, this paper presents PortraitTalk, a robust and customizable audio-to-talking face generation framework that integrates audio, visual, and textual modalities to enable fine-grained facial attribute control, coherent background generation, and accurate audio–lip synchronization for unseen identities. Specifically, PortraitTalk comprises two pivotal components: IdentityNet and AnimateNet.
IdentityNet preserves consistent identity across the generated video frames, while integrating semantics from text descriptions into the generated faces. To enable customization, we integrate text and image embeddings in IdentityNet through the decoupled cross-attention mechanism [21]. Traditional approaches typically concatenate text-image pairs in text-driven talking face generation [18, 22] or audio-image pairs in audio-driven talking face generation [23, 13, 24, 25] within a singular attention block. We argue that this approach restricts a generative model from thoroughly learning and integrating identity-specific facial details and textual descriptions. To further enhance generation quality, we incorporate a mask reconstruction loss into IdentityNet. In this setup, a portion of input frames is randomly corrupted, and the model is trained to reconstruct the original frames. Inspired by masked language modeling [26] in natural language processing, this strategy encourages the model to infer missing visual information from the surrounding context and guiding conditions. The underlying hypothesis is that by learning to recover corrupted regions using unmasked areas, the network implicitly gains a stronger understanding of visual structure and facial coherence, leading to improved generation fidelity. Through extensive experiments, we find that IdentityNet, enhanced with mask reconstruction and decoupled cross-attention, not only preserves high-fidelity identity across frames but also enables consistent talking-face generation without altering core facial features. By integrating textual semantic description within the latent space, it supports fine-grained editing, such as adjusting hairstyle, hair color, background, or environmental settings, all without the need for post-processing.
Other critical quality factors in talking face generation are audio-lip synchronization and temporal consistency. Even minor artifacts or inconsistent head movements can noticeably reduce visual realism. To address this, AnimateNet goes beyond simply mapping audio to facial motion. It leverages a motion generator [27] to extract facial motion representations from audio, forming the first phase of the animation pipeline. This stage improves the generalization capability of PortraitTalk to diverse speech beyond the training data. In the second phase, AnimateNet transforms these motion cues into realistic video frames through a diffusion-based architecture enhanced with dedicated cross-attention mechanisms. Specifically, it incorporates structure-aware attention for precise audio-lip synchronization, identity-guided attention for consistent head placement, and temporal attention for smooth transitions between frames. This design ensures accurate lip movements, consistent head placement, and smooth transitions, resulting in coherent and identity-preserving talking face videos.
@@ I Introduction
Despite the notable progress in audio-to-talking face generation, the existing evaluation metrics often concentrate solely on isolated attributes such as visual quality or lip synchronization. Consequently, these metrics cannot provide a comprehensive assessment that reflects the holistic performance of a method. There exists an imperative need for a new metric that encompasses both spatial and temporal consistencies to facilitate thorough evaluation. Spatial consistency is crucial as it ensures the coherence and stability of the generated video frames, thereby preserving the integrity of identity and expression throughout the video. Temporal consistency, on the other hand, assesses the smoothness and natural flow of the transition over time, which are vital for producing lifelike and persuasive talking faces. In this paper, we propose a novel Audio-Driven Facial Dynamics (ADFD) measurement that addresses the above deficiencies by integrating both spatial and temporal factors, thereby providing a reliable measure of the overall quality of a generated talking face video.
Figure 1: Talking faces generated by PortraitTalk. Given the reference images of an identity and the corresponding audio input, PortraitTalk synthesizes high-quality talking face videos that closely preserve the identity’s appearance and speaking style. Furthermore, the visual attributes of the generated video, such as hair color, age, environmental settings and facial expressions, can be flexibly customized using a simple text prompt. This enables fine-grained control over the generated content, allowing users to tailor the visual presentation to specific narrative or stylistic requirements.
By tackling the aforementioned challenges, PortraitTalk represents a significant step forward in audio-to-talking face generation, offering a more robust and flexible solution that enhances the realism and practicality of generated talking face videos. Some examples of the generated talking faces by PortraitTalk are shown in Fig. 1. Separating IdentityNet and AnimateNet allows each module to specialize in distinct aspects of the generation process. IdentityNet focuses on preserving the identity and reflecting text prompts, while AnimateNet is designed to handle motion-related information by mapping audio-driven motion priors and applying geometry control conditions to the vision domain. In summary, the main contributions of PortraitTalk include:
@@
  • We develop a robust and customizable audio-to-talking face generation framework that can generalize well to unseen faces. We use three different modalities in the proposed framework.
@@
  • We propose a new evaluation metric that effectively measures video quality in terms of both spatial and temporal aspects.
@@
  • Qualitative and quantitative experiments demonstrate the effectiveness of PortraitTalk in enhancing synchronization, realism, and customization of generated talking faces as compared with the state-of-the-art methods.
The rest of this paper is organized as follows: We first introduce the existing literature in Section II. Then Section III presents the proposed PortraitTalk method in detail. Last, the experimental results are reported and analyzed in Section IV, and the conclusion is drawn in Section V.
II Related Work
II-A Audio-driven Talking Face Generation
@@
In the field, various methods have been proposed, with intermediately guided talking face generation being particularly common. Such models leverage intermediate representations, such as 2D facial landmarks or 3D faces, to effectively bridge the gap between audio input and visual output. These methods [5, 28, 29, 30, 11, 31, 32] typically consist of two main components: the first predicts intermediate representations from audio, and the second uses these features to generate talking faces. This two-stage methodology facilitates control over the synchronization between audio and visual data, thereby enhancing the realism and accuracy of the generated talking faces.
@@
Chen et al. [4] proposed a method that converts audio into mouth landmarks to generate talking faces. However, as it only predicts movements for the lower half of the face and uses the upper half from ground truth frames, the resulting videos often lack realism. MakeItTalk [5] introduced a one-shot audio-to-talking-face generation model based on facial landmarks. While it can handle various identities, MakeItTalk struggles with precise audio-lip synchronization and often fails to deliver high-quality talking faces. The Audio2Head [33] model generates talking faces by extracting relatively dense motion fields from audio, but the faces often exhibit distortions and lack consistency in identity preservation. SadTalker [34] has significantly enhanced the generalization capability through a learned latent space. Despite this advancement, the head movements, which are generated based on predefined motion coefficients, frequently lack realism. IP-LAP [35], a transformer-based generative model, utilizes audio inputs and sketches to produce talking faces. Although it achieves stable head movements, it falls short in synchronizing lip movements with audio, particularly with identities unseen during training.
Figure 2: PortraitTalk has two main components: IdentityNet and AnimateNet. Text and identity embeddings are derived from the text and face encoder, with a projection layer mapping identity features into the text embedding dimension. These features are integrated into IdentityNet using a decoupled cross-attention mechanism to capture subtle facial characteristics. Simultaneously, facial motions corresponding to the input speech, enhanced by head placement guidance, are processed through AnimateNet to ensure dynamic and temporal coherence. In PortraitTalk, a latent diffusion model serves as the foundational rendering mechanism. The structural attention block incorporates head placement guidance and facial landmark mapping. For simplicity, these elements are represented within a single block.
Recently, Diffusion Models (DM [36]) have been increasingly adopted for talking face generation [37, 38, 39, 40, 41], offering improvements in both visual quality and audio-lip synchronization. Despite these advances, such models often struggle with generalization, particularly when presented with identities that differ significantly from those seen during training. Some approaches [38, 39], exhibit limited audio-lip synchronization, which undermines the naturalness of the output. Furthermore, fine-grained control and customization in audio-driven talking face generation remain underexplored in existing diffusion-based models.
II-B Emotional Talking Face Generation
@@
Emotional talking face generation has attracted increasing attention, primarily due to its potential to enhance communication and its utility in the entertainment industry. Traditional methods have predominantly used discrete emotion labels to model expressions. For example, [42] developed a model inspired by Wav2Lip [12], integrating an emotion label encoder, emotion-specific loss function, and an emotion discriminator. This model, however, primarily modifies the mouth region, thereby producing videos that lack realistic appearance variations. Furthermore, it exhibits limited effectiveness in displaying diverse expressions and fails to maintain an identity consistency across the generated frames.
@@
Label-based emotional talking face generation typically struggles to produce controllable and fine expressions. To address these limitations, recent developments advocate the use of emotional videos as references [43, 44, 13, 2, 45]. This approach allows for the injection of expressions directly derived from other videos, enhancing the expressiveness of generated faces. For example, EAMM [13] uses an emotional source video to generate emotional talking faces. Nonetheless, this method often struggles with identity consistency, exhibits noticeable irregularities, and encounters audio-lip synchronization issues. These challenges underscore the necessity for the development of a model that can integrate emotion with audio-visual semantics more effectively. Similarly, EDTalk [2] employs emotions from reference-style videos to enhance the expressiveness of talking faces. While this method substantially enhances expressiveness, it sometimes does so at the cost of visual quality.
III The Proposed Method
@@
PortraitTalk is a customizable audio-driven talking face generation framework that synthesizes identity-preserving facial animations with accurate audio-lip synchronization, using reference images and the corresponding speech. Fig. 2 illustrates the pipeline of PortraitTalk. To enhance customization and controllability, we incorporate text prompt embeddings, enabling the model to edit facial attributes such as style, expression, and appearance during the generation process. The framework consists of two main components. The first, IdentityNet, extracts both high-level semantic information and low-level visual details from the reference image and text prompt, ensuring consistent identity preservation across frames. The second, AnimateNet, conditions the generation on speech-driven motion features and employs temporal mechanisms to maintain frame-to-frame coherence and enable smooth, realistic transitions throughout the video.
III-A IdentityNet
IdentityNet is built upon the architecture of Stable Diffusion [46] (SD 1.5) and serves as the primary component for preserving facial identity throughout the generated video. The backbone is initialized with the pre-trained UNet weights from SD 1.5, enabling the model to retain the strong generative capacity of the original network while adapting to the talking face generation task through fine-tuning on our dataset. This initialization provides a stable foundation for IdentityNet.
To improve the consistency of synthesized frames and enhance the reconstruction capability of IdentityNet, we apply a masked fine-tuning strategy, as depicted in Fig. 3. Specifically, random regions of the input frames are corrupted with Gaussian noise, and the model learns to reconstruct the missing content. Unlike conventional zero-masking approaches, the use of Gaussian noise aligns more naturally with the denoising objective of diffusion models. This strategy encourages the model to attend to global facial structure, such as head shape and overall appearance, rather than overfitting to local pixel patterns, thereby promoting more stable and identity-consistent generations [47].
@@ III-A IdentityNet
The fine-tuning process follows the training strategy of stable diffusion. Randomly masked frames are encoded by passing through the encoder to obtain a masked latent representation $\mathbf{z}_{m}$. Subsequently, a forward and backward diffusion process is applied across $T$ time steps within this masked latent space. The denoising U-Net ($\epsilon_{\theta}$) is trained to reconstruct the original image by estimating the amount of noise ($\epsilon$) introduced at each forward step. This is achieved by the mask reconstruction loss:
@@
$L^{mask}=\mathbb{E}_{t,\mathbf{z}_{t},c,\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{m_{t}},t,\mathbf{c})\|^{2}\right]$ (1)
@@
where $t$ is a sampled time step, ${z}_{m_{t}}$ is the noisy masked latent at time step $t$, and $c$ is the condition set (text and reference images). This loss function guides the network towards accurate noise estimation and effective image restoration.
Figure 3: An overview of the masked loss fine-tuning strategy used for IdentityNet. During training, random regions of the input frames are corrupted, and the model is optimized to reconstruct the original content. This masked fine-tuning approach encourages the network to focus on the global facial structure and identity-relevant features, rather than overfitting to local pixel-level details. The process operates in the latent space of a diffusion model: the masked input is first encoded, followed by a forward and backward diffusion process over multiple time steps ($T_{steps}$). The denoising U-Net is trained to estimate the added noise using a mask reconstruction loss ($L_{mask}$), which guides IdentityNet toward producing more consistent, stable, and identity-preserving generations.
@@
Despite the effectiveness of SD in image generation, it struggles to maintain consistent identity characteristics across consecutive frames. Moreover, given our model’s goal of enabling customization during the generation process, it is essential to incorporate user prompts, both textual and visual, into the denoising blocks effectively. This ensures that the generated talking faces consistently exhibit facial details that align with the input text and reference image. To tackle these challenges, we employ a decoupled cross-attention mechanism that processes text and image features independently, enabling more targeted and coherent feature integration. Specifically, we apply the decoupled cross-attention module [21] to process text and reference image features independently, embedding semantic representations from both modalities into the IdentityNet backbone to guide the generation process.
@@
In preceding studies [21, 48], CLIP has been widely used as an image encoder. However, it has proven inadequate for extracting strong facial features in audio-to-talking face generation. This deficiency often results in unstable facial representations across frames, while barely reflecting text prompt guidance in generated videos. CLIP, a contrastive language–image model, is primarily trained on text-image pairs from the natural image domain. This training predisposes CLIP to prioritize broad and ambiguous features such as color and style, which are less effective for detailed face image understanding [49]. Such limitations pose significant challenges for the proposed framework. This task necessitates precise identity preservation, wherein robust semantic representation of facial features is crucial. To achieve this, we initially eliminate background elements from reference images and subsequently extract facial embeddings utilizing a face recognition model [49]. This approach has demonstrated enhanced facial fidelity across generated frames. Simultaneously, text prompts are processed using the CLIP text encoder, and facial embeddings are mapped to the same dimensional space as text embeddings via a projection layer. Both text and facial representations are then integrated into the denoising UNet using decoupled cross-attention blocks, enhancing the facial fidelity and editability of the generated talking faces.